# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ Acknowledgements
This work is supported in part by the Arizona Technology and Research Initiative Fund (TRIF) for Innovative Technologies for the Fourth Industrial Revolution.
@@
In this section, we introduce a collection of technical lemmas necessary to support the results presented in the paper.
8 Required Lemmas
@@
$\leq\frac{\beta}{1-\beta}\sum_{k=0}^{K-1}A_{k}^{2}.$ (74)
@@
$\quad+C_{\theta}^{\Phi}\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}-[\nabla_{\theta\theta}^{2}g(\overline{x},\theta^{*}(\overline{x}))]^{-1}\right\|.$ (76)
where in the last inequality we used Assumptions 2.1 and 2.2-(iii) along with the premises of Assumption 2.3. Moreover, for any invertible matrices $H_{1}$ and $H_{2}$, we have that
@@
Therefore, using the result of Lemma 2.2-(I) and (77) we can further bound inequality (8.1) as follows,
@@
$\leq(\frac{L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu_{g}}+\frac{C_{\theta}^{\Phi}L_{\theta\theta}^{g}}{\mu^{2}_{g}}(1+\mathbf{L}_{\theta}))\left\|x-\overline{x}\right\|+\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}\left\|\overline{y}-y\right\|.$ (78)
@@
Note that $\nabla_{\theta}g(x,\theta^{*}(x))=\nabla_{\theta}g(\overline{x},\theta^{*}(\overline{x}))=0$. Using the Cauchy-Schwartz inequality we have
@@
$\leq C_{\theta x}^{g}\left\|x-\overline{x}\right\|\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|.$
where the last inequality is obtained by using Assumption 2.2. Therefore, we conclude that $\mu_{g}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|\leq C_{\theta x}^{g}\left\|x-\overline{x}\right\|$ which leads to the desired result in part (I).
Part (II): We start proving this part using the definition of $\nabla_{x}\mathcal{L}(x,y)$ stated in (8a). Using the triangle inequality we obtain,
@@
$\quad+\left.[\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(x,y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))v(x,y)]\right\|,$ (79)
@@
Next, from Assumptions 2.1-(1) and 2.2-(ii) together with the triangle inequality we conclude that
@@
$\quad+C_{\theta x}^{g}\left\|v(\overline{x},\overline{y})-v(x,y)\right\|+\frac{C_{\theta}^{\Phi}}{\mu_{g}}\left\|\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))\right\|$ (80)
@@
Combining the result of Lemma 2.2-(I) and Assumption 2.3 with the Assumption 2.2-(iv) leads to
@@
$\quad+(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})\left\|y-\overline{y}\right\|,$ (81)
@@
$\leq(L_{yx}^{\Phi}+L_{y\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|.$ (82)

### v3

@@ Acknowledgements
This work is supported in part by NSF Grant 2127696 and the Arizona Technology and Research Initiative Fund (TRIF) for Innovative Technologies for the Fourth Industrial Revolution.
@@
8 Auxiliary Lemmas
@@
$\leq\frac{\beta}{1-\beta}\sum_{k=0}^{K-1}A_{k}^{2}.$
@@
$\quad+C_{\theta}^{\Phi}\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}-[\nabla_{\theta\theta}^{2}g(\overline{x},\theta^{*}(\overline{x}))]^{-1}\right\|,$ (34)
where in the last inequality, we used Assumptions 2.1 and 2.2-(3) along with the premises of Assumption 2.3. Moreover, for any invertible matrices $H_{1}$ and $H_{2}$, we have that
@@
Therefore, using the result of Lemma 2.2-(I) and (35), we can further bound inequality (8.1) as follows,
@@
$\leq(\frac{L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu_{g}}+\frac{C_{\theta}^{\Phi}L_{\theta\theta}^{g}}{\mu^{2}_{g}}(1+\mathbf{L}_{\theta}))\left\|x-\overline{x}\right\|+\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}\left\|\overline{y}-y\right\|.$
@@
Note that $\nabla_{\theta}g(x,\theta^{*}(x))=\nabla_{\theta}g(\overline{x},\theta^{*}(\overline{x}))=0$. Using the Cauchy-Schwartz inequality, we have
@@
$\leq C_{\theta x}^{g}\left\|x-\overline{x}\right\|\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|,$
where the last inequality is obtained by using Assumption 2.2. Thus, we conclude that $\mu_{g}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|\leq C_{\theta x}^{g}\left\|x-\overline{x}\right\|$ which leads to the desired result in part (I).
Part (II): We start proving this part using the definition of $\nabla_{x}\mathcal{L}(x,y)$ stated in (4a). Using the triangle inequality, we obtain,
@@
$\quad+\left.[\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(x,y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))v(x,y)]\right\|,$
@@
Next, from Assumptions 2.1-(1) and 2.2-(2) together with the triangle inequality, we conclude that
@@
$\quad+C_{\theta x}^{g}\left\|v(\overline{x},\overline{y})-v(x,y)\right\|+\frac{C_{\theta}^{\Phi}}{\mu_{g}}\left\|\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))\right\|.$
@@
Combining the result of Lemma 2.2-(I) and Assumption 2.3 with the Assumption 2.2-(4) leads to
@@
$\quad+(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})\left\|y-\overline{y}\right\|,$
@@
$\leq(L_{yx}^{\Phi}+L_{y\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|,$